HUBBLE SPACE TELESCOPE DAILY REPORT #3001 PERIOD COVERED: 0000Z (UTC) 11/19/01 - 0000Z (UTC) 11/20/01 Daily Status Report as of 324/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Six Sets of WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.2 Completed WF/PC-2 9227 (Observations of Astrophysically Important Visual Binaries) The WF/PC-2 was used to perform observations of three visual binary stars that will ultimately yield fundamental astrophysical results, once their orbits and masses are determined. The targets are the following: {1} Procyon {P=41 yr}, for which our first WF/PC-2 images yielded an extremely accurate angular separation of the bright F star and its very faint white-dwarf companion. Combined with ground-based astrometry of the bright star, our observation significantly revised downward the derived masses, and brought Procyon A into excellent agreement with theoretical evolutionary tracks for the first time. The mass of Procyon B, however, implies a completely unexpected chemical composition for the white dwarf, and now poses a sharp evolutionary puzzle. With the continued monitoring proposed here, we will obtain masses to an accuracy of better than one percent, providing a test bed for theories of Sun-like stars and white dwarfs. {2} G 107-70, a close double white dwarf {P=19 yr} that promises to add two accurate masses to the tiny handful of white-dwarf masses that are directly known from dynamical measurements. {3} Mu Cas {P=21 yr}, a famous metal-deficient G dwarf for which accurate masses will lead to the stars' helium contents, with cosmological implications. No problems were encountered for these observations. 1.3 Completed Three Sets of WF/PC-2 8942 (Cycle 10 Intflat Sweeps and Linearity Test) The WF/PC-2 was used to monitor the pixel-to-pixel flatfield response and provide a linearity check. These intflat sequences are done once during the year and the images will provide a backup database in the event of complete failure of the visflat lamp as well as allow monitoring of the gain ratios. There were no reported problems. 1.4 Completed Three Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.5 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 1.6 Completed WF/PC-2 8726 (Cometary Knots in Planetary Nebulae) The WF/PC-2 was used to make second epoch observations of the Cometary Knots in NGC 8653 in order to get a better idea of just how these objects are formed by comparison of their spatial motions with that of the ambient gas. The observations completed with no reports of problems. 1.7 Completed Two Sets of STIS/CCD 8906 (Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to measure the effectiveness of the CCD hot pixel annealing process by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. In addition, CTE performance is examined by looking for traps in a low signal level flat. There were no reported problems. 1.8 Completed WF/PC-2 9091 (Evolution of Star Forming Environments in H II Regions) The WF/PC-2 was used to look for evaporating gaseous globules and silhouette disks in two young clusters and along one molecular cloud wall. The proposal completed nominally. 1.9 Completed STIS/CCD 8611 (Observations of Nearby Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to obtain UV spectra of five nearby {0.02 < z < 0.08} SNe Ia in the Hubble Flow. The spectra will be taken at weekly intervals over a range in time starting slightly before maximum light and extending to +30 days. These observations will accomplish the following three goals: {1} calibration of the rest frame UV light curves of SNe Ia and an assessment of their potential use as distance indicators through UV light curve shape analyses. {2} improvement in our understanding of the physics of SNe Ia, metallicity/evolutionary effects and correlations between peak brightness and UV spectral features. {3} calibration of the SNe Ia previously observed by HST at high-redshift. This data is crucial for proper cross-filter k-corrections and calibration of the supernova photometry. The observations completed with no reported anomalous activity. 1.10 Completed Three Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.11 Completed STIS/CCD 9106 (The Biggest Black Holes) The Space Telescope Imaging Spectrograph (CCD) was used to perform searches for supermassive black holes in galaxy centers that have led to the discoveries that {1} most or all hot galaxies contain massive dark objects at their centers, presumably black holes; and {2} there is a tight correlation between the black-hole mass and the luminosity-weighted velocity dispersion of the hot component of the galaxy. This remarkable relationship suggests a strong link between black-hole formation, AGN activity, and galaxy formation, and once it is understood this link should advance our understanding of all three processes. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 11 Successful: 11 2.2 FHST Updates: Scheduled: 16 Successful: 16 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared five times. The engineering status buffer was dumped and the limits reset (SESBSLD) at 323/1449Z per ROP DR-18A. The ephemeris table was uplinked at 324/0001Z per ROP DF-01A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW